OpenStudy (anonymous):

will give out medals. Which fraction is equivalent to 0.2?

OpenStudy (anonymous):

1/5, 2/10, 3/15, 4/20

OpenStudy (anonymous):

0.2 = 2/10 --> in lowest terms would be 1/5 so 1/5 would be your answer

OpenStudy (anonymous):

2/10 = 1/5